# Vulnerability-Aware Spatio-Temporal Learning for Generalizable and Interpretable Deepfake Video Detection

### v2

@@ 5 Conclusion
This paper introduces a fine-grained approach for generalizable and interpretable deepfake video detection with two main contributions. First, we propose a multi-task learning framework that addresses both spatial and fine-grained temporal vulnerabilities in high-fidelity deepfake videos, incorporating a standard classification branch along with two new auxiliary branches (temporal and spatial). These proposed branches help the model focus on vulnerable regions and provide interpretable insights into how the network ”sees” the data. Second, we present a high-quality pseudo-fake generation technique to improve model generalization by creating rich, generic spatio-temporal artifacts. Extensive experiments on challenging benchmarks demonstrate that FakeSTormer achieves competitive performance compared to SOTA methods while providing interpretability. Future work will investigate pretraining strategies for network initialization.
@@ 6 Appendix
6.1 Code Release
The code is provided as a part of the supplementary materials. The code is under license registration and will be publicly released with pretrained weights.
@@ 6.2 Details of the Ground-Truth Derivative
Normalization functions. Employing a normalization function is important for stabilizing the training of our temporal branch $h$. Therefore, we consider three different normalization functions including Standardization (MeanStd), MinMax, and Unnormalized 3D Gaussian. Specifically, for the Standardization and MinMax, we respectively compute the std($\mathbf{D}$)-mean($\mathbf{D}$), and min($\mathbf{D}$)-max($\mathbf{D}$), while we follow the work of [41] to adapt an unnormalized Gaussian map from 2D to 3D for normalization. We report in Table. 8 cross-evaluation results on five datasets [35, 16, 13, 14, 68] with the use of the three investigated functions, using a model trained on FF++ [45]. It can be noted that the model is robust to various types of normalization functions with the best performance recorded for the Standardization approach.
@@
CDF2 DFD DFDCP DFDC DFW Avg.
MinMax 91.6 95.9 88.1 71.1 70.1 83.4
Unnormalized 3D Gaussian 92.5 91.8 86.4 72.4 76.3 83.9
MeanStd 92.4 98.5 90.0 74.6 74.2 85.9
@@
6.3 SBVs: Pseudo-code and Visual Samples
Algorithm. To enhance the clarity and reproducibility of the SBVs generation process, we provide the overall algorithm in the form of pseudo-code, as detailed in Algorithm 1.
Figure 7: Illustration of the facial landmarks, the generated SBV, and the blending boundary with and without applying the Consistent Synthesized Parameters (CSP) module and the Landmark Interpolation module.
@@ 16             $\mathbf{M}_{i}\leftarrow\mathbf{M}_{i}\cup\{\mathbf{M}_{i}(t)\}$
Algorithm 1 Pseudo-code for SBVs Generation
@@
Visual Samples. To visually demonstrate the benefits of the Consistent Synthesized Parameters (CSP) and the Landmark Interpolation module (Section 3.1 in the main paper) in generating high-quality pseudo-fake videos, we show some SBV samples, their blending boundaries, original landmarks, and those modified by the proposed modules in Figure 7. In the top part of the figure, we compare data generated using only CSP to data generated with both CSP and the Landmark Interpolation module. We observe that the Landmark Interpolation module ensures smooth transitions of facial landmarks between consecutive frames ($t\rightarrow t+1$). In the bottom part of the figure, we compare data generated with only CSP to data generated without any of the proposed SBVs components. We observe significant variations in the manipulated facial areas when CSP is omitted. Therefore, the proposed CSP and Landmark Interpolation module effectively enhance the temporal coherence of the generated SBVs.
6.4 Impact of SBVs
@@
To verify the advantage of using SBVs for improving the generalization of detectors, we conduct several experiments using different binary classifiers trained either with SBVs or with one of the four types of forgeries forming FF++ [45] (DF [10], F2F [50], FS [30], NT [49]). For a fair comparison, a widely-used CNN-based Resnet3D [26] and a Transformer-based TimeSformer [2] are employed. We note that both selected models are trained from Scratch (S) without pretrained initialization. Table. 9 presents the generalization performance in terms of AUC (%) on five datasets [35, 16, 13, 14, 68] respectively when trained with different manipulation methods from FF+. Notably, training with SBVs significantly increases the overall generalizability capability of binary models as compared to those trained on using one specific manipulation. This indicates the importance of highly realistic, naturally consistent generated pseudo-fake videos.
Method Training set Test set AUC (%)
Real DF FS F2F NT FF++ CDF2 DFD DFDCP DFDC DFW Avg.
ResNet3D [26] $\times$ $\times$ $\times$ 72.5 58.5 51.3 53.4 59.4 65.0 60.0
TimeSFormer [2] $\times$ $\times$ $\times$ 65.4 59.3 66.1 53.5 61.4 57.5 60.5
ResNet3D [26] $\times$ $\times$ $\times$ 70.6 61.1 50.6 59.2 55.8 51.5 58.1
TimeSFormer [2] $\times$ $\times$ $\times$ 76.4 51.7 43.7 44.6 54.5 43.9 52.5
ResNet3D [26] $\times$ $\times$ $\times$ 78.0 63.8 54.5 63.4 55.7 50.1 60.9
TimeSFormer [2] $\times$ $\times$ $\times$ 81.1 64.4 60.1 64.5 52.0 50.5 62.1
ResNet3D [26] $\times$ $\times$ $\times$ 72.7 63.7 75.6 69.1 59.6 62.6 67.2
TimeSFormer [2] $\times$ $\times$ $\times$ 75.5 65.8 84.7 70.3 62.7 65.5 70.8
ResNet3D [26] + SBVs $\times$ $\times$ $\times$ $\times$ 90.2 85.9 85.0 82.8 66.4 67.5 79.6($\uparrow$12.4)
TimeSFormer [2] + SBVs $\times$ $\times$ $\times$ $\times$ 94.7 89.5 95.6 88.6 72.5 70.9 85.3($\uparrow$14.5)
@@
Table 9: Cross-dataset generalization. Performance comparison in terms of AUC (%) on multiple datasets of different binary classification models [2, 26] trained using our video synthesis (SBVs) and normal fake data [45]. All models are trained on FF++(c23) [45] from Scratch (S) and are tested on other datasets [35, 16, 13, 14, 68]. Gray indicates the use of normal fake data for training. Bold and underline highlight the best and the second-best performance, respectively.
@@ 6.5 Robustness to Unseen Perturbations
In the main manuscript, we report in Figure. 5 the “Average” performance under different corruptions. This section complements this experiment by reporting the mean performance across different severity levels for each degradation type, as detailed in Table. 10. Except for a slight decrease in effectiveness under “Color Saturation” compared to LAA-Net [42], FakeSTormer is generally more robust to the unseen perturbations as compared to other augmented-based methods [42, 47, 32, 34].
Method Fake Contrast Saturation Gaussian Blur Gaussian Noise JPEG Compression Block Wise Avg
DSP-FWA [34] $\checkmark$ 80.7 79.6 67.3 61.8 68.0 76.6 72.3
FaceXray [32] $\checkmark$ 88.9 96.0 70.0 58.0 62.2 94.7 78.3
SBI [47] 92.3 92.0 72.7 62.2 79.1 92.2 81.7
LAA-Net [42] 95.0 97.0 73.2 57.5 75.2 94.9 82.1
Ours 97.6 96.3 81.6 65.4 92.9 96.8 88.4
@@ 6.6 Multi-shot Inferences
Table 11: Multi-shot inferences. AUC (%) comparison of our model using different numbers of shots in the cross-dataset setup. The AUC slightly increases with a higher number of shots.
6.7 Visualization of Saliency Maps
6.8 STG [33]: Shuffled Frames can produce obvious Temporal Inconsistencies
We propose SBVs to generate subtler artifacts for pseudo-fakes compared to the STG approach used in [33]. We believe that STG may produce obvious (low-quality) temporal artifacts, as it shuffles frames in the temporal domain, leading to significant inconsistencies. Figure 8 illustrates how shuffling creates noticeable discrepancies between frames. In contrast, our SBVs leverage consecutive frames to produce subtler temporal artifacts while simulating these artifacts in a different manner (as detailed in Section 3.1 of the main manuscript).
@@ 6.9 Details on the Datasets
Datasets. For our experiments, we select datasets that haven typically used as benchmarks in previous works [55, 65, 56, 62, 9, 58, 54, 3]. For both training and validation, we employ FaceForensics++(FF++) [45], which consists of 1,000 real videos and 4,000 fake videos generated using four manipulation methods: (Deepfakes (DF) [10], FaceSwap (FS) [30], Face2Face (F2F) [50], and NeuralTextures (NT) [49]). It can be noted that, for training, we use only the real videos and generate pseudo-fake data using our synthesized method, SBVs. By default, the c23 version of FF++ is adopted, following the recent literature [65, 55, 62, 56, 9].
For further validation, we also evaluate on the following datasets: (1) Celeb-DFv2 (CDF) [35], a well-known benchmark with high-quality deepfakes; (2) DeepfakeDetection (DFD) [16], which includes 3,000 forged videos featuring 28 actors in various scenes; (3) Deepfake Detection Challenge Preview (DFDCP) [13] and (4) Deepfake Detection Challenge (DFDC) [14], a large-scale dataset containing numerous distorted videos with issues such as compression and noise; and (5) WildDeepfake (DFW) [68], a dataset fully sourced from the internet, without prior knowledge of manipulation methods.

### v3

@@ 5 Conclusion
This paper introduces a fine-grained approach for generalizable deepfake video detection with two main contributions. First, we propose a multi-task learning framework that targets both subtle spatial and fine-grained temporal vulnerabilities in high-fidelity deepfake videos, incorporating a standard classification branch along with two new auxiliary branches (temporal and spatial). These proposed branches help the model focus on vulnerable regions and provide more valuable insights into how the network sees the data while offering more robustness to high-quality deepfakes. This framework is further supported by the introduction of a high-quality pseudo-fake generation technique. Extensive experiments on several challenging benchmarks demonstrate that FakeSTormer achieves superior performance compared to SOTA methods.
Acknowledgment
This work is supported by the Luxembourg National Research Fund, under the BRIDGES2021/IS/16353350/FaKeDeTeR, and by POST Luxembourg. Experiments were performed on the Luxembourg national supercomputer MeluXina. The authors gratefully acknowledge the LuxProvide teams for their expert support.
Supplementary Material
@@ 6.1 Details of the Ground-Truth Derivative
Normalization functions. Employing a normalization function is important for stabilizing the training of our temporal branch $h$. Therefore, we consider three different normalization functions including Standardization (MeanStd), MinMax, and Unnormalized 3D Gaussian. Specifically, for the Standardization and MinMax, we respectively compute the std($\mathbf{D}$)-mean($\mathbf{D}$), and min($\mathbf{D}$)-max($\mathbf{D}$), while we follow the work of [41] to adapt an unnormalized Gaussian map from 2D to 3D for normalization. We report in Table. 8 cross-evaluation results on six datasets [36, 17, 13, 14, 74, 70] with the use of the three investigated functions, using a model trained on FF++ [46]. It can be noted that the model is robust to various types of normalization functions with the best performance recorded for the Standardization approach.
@@
CDF DFD DFDCP DFDC DFW DiffSwap Avg.
MinMax 91.6 95.9 88.1 71.1 70.1 92.1 84.8
Unnormalized 3D Gaussian 92.5 91.8 86.4 72.4 76.3 91.7 85.2
MeanStd 92.4 98.5 90.0 74.6 74.2 96.9 87.8
@@
6.2 SBV: Pseudo-code and Visual Samples
Algorithm. To enhance the clarity and reproducibility of the SBV generation process, we provide the overall algorithm in the form of pseudo-code, as detailed in Algorithm 1.
Figure 5: Illustration of the facial landmarks, the generated SBV, and the blending boundaries with and without applying the Consistent Synthesized Parameters (CSP) module and the Landmark Interpolation (LI) module. The lack of applied CSP and LI indicates simply stacked SBIs (BottomRight).
@@ 16             $\mathbf{M}_{i}\leftarrow\mathbf{M}_{i}\cup\{\mathbf{M}_{i}(t)\}$
Algorithm 1 Pseudo-code for SBV Generation
@@
Visual Samples. To visually demonstrate the benefits of the Consistent Synthesized Parameters (CSP) and the Landmark Interpolation (LI) module (Section 3.1 in the main paper) in generating high-quality pseudo-fake videos, we show some SBV samples, their blending boundaries, original landmarks, and those modified by the proposed modules in Figure 5. In the top part of the figure, we compare data generated using only CSP to data generated with both CSP and the Landmark Interpolation module. We observe that the Landmark Interpolation module ensures smooth transitions of facial landmarks between consecutive frames ($t\rightarrow t+1$). In the bottom part of the figure, we compare data generated with only CSP to data generated without any of the proposed SBV components. We observe significant variations in the manipulated facial areas when CSP is omitted. Therefore, the proposed CSP and Landmark Interpolation module effectively enhances the temporal coherence of the generated SBV.
6.3 Impact of SBV
@@
To verify the advantage of using SBV for improving the generalization of detectors, we conduct several experiments using different binary classifiers trained either with SBV or with one of the four types of forgeries forming FF++ [46] (DF [10], F2F [51], FS [31], NT [50]). For a fair comparison, a widely-used CNN-based Resnet3D [27] and a Transformer-based TimeSformer [2] are employed. We note that both selected models are trained from Scratch (S) without pretrained initialization. Table. 9 presents the generalization performance in terms of AUC (%) on five datasets [36, 17, 13, 14, 74] respectively when trained with different manipulation methods from FF+. Notably, training with SBV significantly increases the overall generalizability capability of binary models as compared to those trained on using one specific manipulation. This indicates the importance of highly realistic, naturally consistent generated pseudo-fake videos.
Method Pretrain Training set Test set AUC (%)
Real DF FS F2F NT FF++ CDF DFD DFDCP DFDC DFW Avg.
ResNet3D [27] $\times$ $\times$ $\times$ $\times$ 72.5 58.5 51.3 53.4 59.4 65.0 60.0
TimeSFormer [2] $\times$ $\times$ $\times$ $\times$ 65.4 59.3 66.1 53.5 61.4 57.5 60.5
ResNet3D [27] $\times$ $\times$ $\times$ $\times$ 70.6 61.1 50.6 59.2 55.8 51.5 58.1
TimeSFormer [2] $\times$ $\times$ $\times$ $\times$ 76.4 51.7 43.7 44.6 54.5 43.9 52.5
ResNet3D [27] $\times$ $\times$ $\times$ $\times$ 78.0 63.8 54.5 63.4 55.7 50.1 60.9
TimeSFormer [2] $\times$ $\times$ $\times$ $\times$ 81.1 64.4 60.1 64.5 52.0 50.5 62.1
ResNet3D [27] $\times$ $\times$ $\times$ $\times$ 72.7 63.7 75.6 69.1 59.6 62.6 67.2
TimeSFormer [2] $\times$ $\times$ $\times$ $\times$ 75.5 65.8 84.7 70.3 62.7 65.5 70.8
ResNet3D [27] + SBV $\times$ $\times$ $\times$ $\times$ $\times$ 90.2 85.9 85.0 82.8 66.4 67.5 79.6($\uparrow$12.4)
TimeSFormer [2] + SBV $\times$ $\times$ $\times$ $\times$ $\times$ 94.7 89.5 95.6 88.6 72.5 70.9 85.3($\uparrow$14.5)
@@
Table 9: Cross-dataset generalization. Performance comparison in terms of AUC (%) on multiple datasets of different binary classification models [2, 27] trained using our video synthesis (SBV) and normal fake data [46]. All models are trained on FF++(c23) [46] from Scratch (S) and are tested on other datasets [36, 17, 13, 14, 74]. Gray indicates the use of normal fake data for training. Bold and underline highlight the best and the second-best performance, respectively.
@@ 6.4 Robustness to Unseen Perturbations
In the main manuscript, we report in Figure. 5 the “Average” performance under different corruptions. This section complements this experiment by reporting the mean performance across different severity levels for each degradation type, as detailed in Table. 10. Except for a slight decrease in effectiveness under “Change Saturation” compared to LAA-Net [42], FakeSTormer is generally more robust to the unseen perturbations as compared to other augmented-based methods [42, 48, 33, 35].
Method Real Fake Contrast Saturation Gaussian Blur Gaussian Noise JPEG Compression Block Wise Avg
DSP-FWA [35] $\checkmark$ $\checkmark$ 80.7 79.6 67.3 61.8 68.0 76.6 72.3
FaceXray [33] $\checkmark$ $\checkmark$ 88.9 96.0 70.0 58.0 62.2 94.7 78.3
SBI [48] $\checkmark$ $\times$ 92.3 92.0 72.7 62.2 79.1 92.2 81.7
LAA-Net [42] $\checkmark$ $\times$ 95.0 97.0 73.2 57.5 75.2 94.9 82.1
Ours $\checkmark$ $\times$ 97.6 96.3 81.6 65.4 92.9 96.8 88.4
@@ 6.5 Multi-shot Inferences
Table 11: Multi-shot inferences. AUC (%) comparison of our model using different numbers of inference shots in the cross-dataset setup. The AUC slightly increases with a higher number of shots.
6.6 Visualization of Auxiliary Branches’ Outputs
In addition to the probability output of the standard classification branch, FakeSTormer can provide more valuable insights from our auxiliary branches that might be conductive to prediction’s post-analyses. Specifically, the spatial and temporal branches output the intensity of the spatial artifacts encoded in each video frame and the vulnerability change over time, respectively. The spatial branch provides frame-level scores, while the temporal branch offers more fine-grained insights. As shown in Figure 7, the spatial outputs (denoted by the numbers in each frame) denote high values for fake data and low values for real data. For the temporal outputs, the heatmaps show the change in vulnerability between the instant frame $t$ and the previous one. It can be observed that it primarily focuses around the blending boundaries. We note that the change between $t-1$ and $t$ is visualized at the $t^{th}$ frame; hence, there is a blank heatmap at the $1^{st}$ frame.
Figure 7: Visualization of Auxiliary Branches’ Outputs. We visualize the additional auxiliary spatial and temporal branches’ outputs on different unseen datasets. As shown, the number on each frame denotes the output of the spatial branch $g$, while the heatmap visualizes the output of the temporal branch $h$.
6.7 STC [34]: Shuffled Frames can produce obvious Temporal Inconsistencies
We propose SBV to generate subtler artifacts for pseudo-fakes compared to the STC approach used in [34]. We believe that STC may produce obvious (low-quality) temporal artifacts, as it shuffles frames in the temporal domain, leading to significant inconsistencies. Figure 6 illustrates how shuffling creates noticeable discrepancies between frames. In contrast, our SBV leverages consecutive frames to produce subtler temporal artifacts while simulating these artifacts in a different manner (as detailed in Section 3.1 of the main manuscript).
@@ 6.8 Details on the Datasets
Datasets. For our experiments, we select datasets that haven typically used as benchmarks in previous works [56, 71, 57, 67, 9, 59, 55, 3]. For both training and validation, we employ FaceForensics++(FF++) [46], which consists of 1,000 real videos and 4,000 fake videos generated using four manipulation methods: (Deepfakes (DF) [10], FaceSwap (FS) [31], Face2Face (F2F) [51], and NeuralTextures (NT) [50]). It can be noted that, for training, we use only the real videos and generate pseudo-fake data using our synthesized method, SBV. By default, the c23 version of FF++ is adopted, following the recent literature [71, 56, 67, 57, 9].
For further validation, we also evaluate on the following datasets: (1) Celeb-DFv2 (CDF) [36], a well-known benchmark with high-quality deepfakes; (2) DeepfakeDetection (DFD) [17], which includes 3,000 forged videos featuring 28 actors in various scenes; (3) Deepfake Detection Challenge Preview (DFDCP) [13] and (4) Deepfake Detection Challenge (DFDC) [14], a large-scale dataset containing numerous distorted videos with issues such as compression and noise; (5) WildDeepfake (DFW) [74], a dataset fully sourced from the internet, without prior knowledge of manipulation methods; (6) DiffSwap generated in the similar protocol as in LFGDIN [63] by using a recent diffusion-based swapping method [70] on $250$ real videos selected from CDF [36]; and (7) DF40 [61], a highly diverse and large-scale dataset comprising $40$ distinct deepfake techniques, enables more comprehensive evaluations for the next generation of deepfake detection.